how do you get the equation for the problem:
The perimeter of a triangle is 76 cm. side A of the triangle is twice as long as side B. side C is 1 cm longer than side A. find the length of each side.

Answer by mananth(16946)

how do you get the equation for the problem:
The perimeter of a triangle is 76 cm. side A of the triangle is twice as long as side B. side C is 1 cm longer than side A. find the length of each side.
..
B=x
A=2x
C=2x+1
..
x+2x+1+2x=76
5x+1=76
-1 -1
5x=75
/5
x=15 cm
....
A=30cm
B=15
C=31cm